# Characterizing Vehicle-Induced Distributed Acoustic Sensing Signals for Accurate Urban Near-Surface Imaging

### v2

@@ 2 Introduction
In urban areas, recent studies have utilized traffic-induced DAS data collected from “dark fiber" to characterize near-surface properties such as S-wave velocity and attenuation (Dou et al., 2017; Yuan et al., 2020; Yang et al., 2022; Lei and Wang, 2024; Ji et al., 2024b). This approach is feasible because most telecommunication optical fibers in urban areas are deployed along roadsides, making traffic the primary seismic source for interferometry. However, roadway vehicles have distinct characteristics, such as vehicle weight, size, and driving speed, that impact the induced surface waves and the inversion of near-surface structures. Directly analyzing surface wave signals from vehicles of varying weights and speeds may result in low-quality near-surface imaging. For instance, we observed that picked fundamental mode dispersion curves from vehicle-induced DAS signals are robust at frequencies above 5 Hz but not at lower frequencies. This is because the dominant traffic type, small/light vehicles, primarily excites surface waves at frequencies above 5 Hz (Yuan, 2023). Therefore, understanding the effects of vehicle characteristics on the induced surface waves is important for improving the quality of near-surface imaging using vehicle-induced DAS signals in urban environments.
@@
In this study, we characterize surface waves generated by moving vehicles of different weights/sizes (sizes are closely correlated with weights) and speeds in the Stanford DAS-2 experiment. Building on our recent work (Yuan et al., 2024), we first use a specialized Kalman filter algorithm (Liu et al., 2023) to track vehicle locations and then classify vehicles by the peak values of their average quasi-static signals into weight categories: light-weight (smaller than the mode value; peak value $\leq$0.57), heavy-weight (laying on the distribution tail; peak value $>$1.2), and mid-weight (in between; 0.57 $<$ peak value $\leq$1.2),. Vehicles are also categorized by their arrival times at DAS into traveling speed categories – slow, mid-speed, and fast – using the mean speed plus or minus one standard deviation as thresholds. Vehicle-induced surface wave windows are then selected using the tracked vehicle trajectories. Virtual shot gathers (VSGs) for the same class of vehicles are retrieved from their surface wave windows and stacked to increase the signal-to-noise ratio (Yuan et al., 2024). A VSG is a shot gathered as if a seismic source were excited at a specific location. This allows us to analyze the phase velocity dispersion at the location and further provides insights into how vehicle characteristics influence the induced surface waves. We also quantify the uncertainties in the picked dispersion curves as the number of stacked VSGs increases by bootstrapping the standard deviation of the phase velocities. Our results show that stacking VSGs reduces the uncertainty of picked dispersion curves, with convergence occurring after signals from approximately 55 vehicles are stacked. Heavy-weight vehicles show higher signal-to-noise ratios and reduced uncertainty in phase velocity estimates, particularly at lower frequencies (e.g., <5 Hz), allowing for better constraints of near-surface structures at greater depths. We also find that differences in the picked dispersion curves for vehicles with different speeds are less pronounced than those due to vehicle weight. Our results suggest that judiciously selecting certain classes of vehicles can be beneficial in processing vehicle-induced surface waves. In particular, for fine-grained temporal monitoring (e.g., daily monitoring) of near-surface characteristics, the low-frequency phase velocities (e.g., $<$5 Hz) should be derived from stacked heavy-weight vehicle VSGs and the high-frequency velocities from stacked VSGs of all classes of vehicles until convergence is achieved; for coarse-grained monitoring (e.g., weekly monitoring), the best dispersion curve should come from stacked heavy-weight vehicle VSGs, which is more efficient and accurate.
@@ 3.1 Surface wave window selection based on vehicle tracking and characterization
Additionally, vehicle weights are estimated using the quasi-static component of the vehicle-induced DAS signals. Assuming linear elasticity of the road and near-surface structures, the prominence amplitude of the quasi-static signal is approximately proportional to the vehicle’s weight. Heavier vehicles produce larger prominence amplitudes. Therefore, vehicles’ relative weights are estimated by computing the average of the quasi-static signal prominence, $\frac{1}{K}\sum_{k=1}^{K}P_{k}$, where $P_{k}$ is the prominence amplitude of the vehicle’s quasi-static signal at DAS channel $k$. The vehicle’s axle length and configuration can also be estimated from its quasi-static signal (Liu et al., 2023; Yuan et al., 2024). Yet, since vehicle length and weight are closely correlated, this study simplifies the analysis without considering different axle lengths and configurations.
@@ 5.1 Characterization of vehicle-induced DAS signal
We begin by characterizing the vehicle-induced DAS signals based on the estimated weights/sizes of the vehicles. To control for the variable of traveling speed, we select vehicles traveling within one standard deviation of the mode speed of all vehicles. We extract the quasi-static signals for each DAS channel by isolating a time window from the band-passed DAS signal (0.1–1 Hz), centered on the vehicle’s arrival time at that channel. Figure 2a displays the histogram of peak values of the average quasi-static signals of these vehicles. We categorize the vehicles into three classes based on the peak values of their average quasi-static signals: vehicles with peak values smaller than the mode (which equals 0.57) are classified as light; those with peak values laying on the distribution tail (peak value $>1.2$) are classified as heavy; and those with values in between these thresholds are considered mid-weight. During the half-month period, we recorded 734 light, 1,058 mid-weight, and 103 heavy vehicles, respectively. Heavy vehicles, on average, weigh approximately seven times more than light vehicles. Figure 2b illustrates the average quasi-static DAS signals for vehicles having different weights, where the shaded areas highlight the signal ranges. Figure 2c shows the power spectral densities (PSD) of the surface wave signals band-passed between 1 and 30 Hz for different vehicle weights, with shaded areas representing the range of PSDs. We observe that heavier vehicles tend to produce both quasi-static and surface wave signals with stronger amplitudes. These signals also exhibit higher signal-to-noise ratio and greater surface wave propagation distances, as evidenced in the example signals (Figures 2d-f). Additionally, surface waves induced by heavy vehicles have a more prominent mode in the low-frequency range (2.5–5 Hz) compared to lighter ones.
@@
We also investigate the DAS signals induced by vehicles traveling at different velocities, as illustrated in Figure 3. Initially, to minimize the influence of vehicle weight and size, we select vehicles whose peak quasi-static signals are within one standard deviation of the mode peak value. Subsequently, these vehicles are categorized into slow, mid-speed, and fast groups based on their average speeds, separated using one standard deviation ($\mu_{s}\pm\sigma_{s}$), as shown in Figure 3a. There are 336 slow, 1,442 mid-speed, and 330 fast vehicles, respectively. Figures 3d–f show the signal examples from vehicles traveling at different speeds with their traveling trajectories indicated by dashed lines, where a smaller slope corresponds to a faster speed. Given the limited range of driving velocities, the surface wave signals among vehicles with varying speeds are not significantly different. Still, we note that surface waves induced by faster vehicles show slightly stronger amplitudes, which could be attributed to the greater kinetic energy and the corresponding larger impact forces exerted on rough road surfaces (Jazar and Marzbani, 2024). Further research could explore the potential for more energetic surface wave signals generated by vehicles traveling at higher velocities, such as those on freeways, where pronounced vehicle-road interactions are likely due to increased velocity and the road roughness (Yuan, 2023).
@@ 5.2 Surface wave dispersion analysis
We find that the quality of both the VSGs and dispersion measurements improves with increasing vehicle weight. The retrieved VSGs from heavier vehicles (Figure 4c) exhibit less noise in the negative time lag section, likely due to more accurate vehicle tracking and reduced ambient noise, as discussed in Yuan et al. (2024). On the dispersion spectra (Figures 4d–f), the fundamental mode of dispersion shows greater continuity as vehicle weight increases, especially at low frequencies (<5 Hz). From the picked dispersion curves for vehicles of different weights (Figures 4g–i), we observe that those from heavy vehicles are more concentrated and show less uncertainty than those from light and mid-weight vehicles. Thus, low-frequency dispersion curves, which resolve deeper structures, can be more reliably measured from heavy vehicles. Retrieving VSGs from heavy vehicles also improves the measurement of higher-order dispersion. Compared to the results from light vehicles (Figure 4d), the higher mode dispersion curves are more distinguishable for mid-weight and heavy vehicles (Figures 4e and f).
@@ 5.3 Inversion of shear-wave velocity
Beyond analyzing the features of the vehicle-induced DAS signals and the dispersion spectra, we inverted the dispersion curves to probe the subsurface structures. We invert for parameters including thickness and shear-wave velocity ($V_{s}$) for each layer, assuming a six-layer model. The density of each layer is approximated using an empirical relationship with the P-wave velocity as in Ji et al. (2024a). From the VSGs, we estimate that the ratio between P-wave and S-wave velocities ranges from 2 to 4 at this location by measuring their travel time, which gives the range of the Poisson ratio between 0.33 and 0.47. The parameter ranges for inversion are specified in Table LABEL:tab:param.
@@ 6.1 Uncertainty quantification of dispersion curves
We gradually increase $N$, the number of VSGs stacked, to study the convergence of the uncertainty of the fundamental mode dispersion curves and to compare the convergence among different vehicle classes. For each bootstrapping, M is set to 30 to ensure sufficient samples for calculating the standard deviation while maintaining manageable data volumes for processing. We apply the bootstrapping method at both pivot locations of 600 and 700 meters. Figures 8a and c show that the uncertainties (standard deviation) converged after stacking approximately 55 VSGs across all cases. The 600-meter pivot location shows a similar result, as detailed in the Supplemental Materials (Figure S3). The converged standard deviation of dispersion curves for heavy vehicles is about four times smaller than that for light and mid-weight vehicles at the 700-meter pivot location and five times smaller at the 600-meter location. The converged standard deviations of dispersion curves for vehicles with different speeds are similar. However, stacking VSGs from fewer slow-moving vehicles (e.g., fewer than 10 vehicles) results in larger uncertainty in dispersion curves.
@@
Figure 8: Uncertainty quantification of the picked dispersion curves for different numbers of vehicles with varying weights and speeds. The uncertainties are quantified by bootstrapping the standard deviations of the estimated fundamental mode dispersion curves. (a) shows the standard deviations of the fundamental mode dispersion curves for vehicles with different weights at pivot locations of 700 meters. (b) displays the converged standard deviations of the fundamental mode dispersion curves at various frequencies for vehicles with different weights. (c) and (d) present the same data as (a) and (b) but for vehicles with different speeds. The color version of this figure is available only in the electronic edition.
@@
Additionally, we investigate the converged uncertainty at different frequencies. Figure 8b and d display the converged standard deviations of the fundamental mode phase velocities at various frequencies for different vehicle classes. At frequencies below 5 Hz, the converged uncertainties for heavy vehicles are significantly smaller than those for light and mid-weight vehicles. For example, at 2.5 Hz, the converged uncertainty for heavy vehicles is approximately 3$\times$ smaller than that for light and mid-weight vehicles at the 700-meter pivot location. Another pivot location at 600 meters presents similar results, as shown in the Supplemental Materials (Figure S3). At higher frequencies, the converged uncertainties of phase velocities for vehicles of different weights are similar. Moreover, the converged uncertainties of phase velocities at different frequencies remain consistent across vehicles traveling at different speeds.

### v3

@@ 2 Introduction
In urban areas, recent studies have utilized ambient noise interferometry to analyze DAS data collected from “dark fiber" to characterize near-surface properties such as S-wave velocity and attenuation (Dou et al., 2017; Yuan et al., 2020; Yang et al., 2022; Lei and Wang, 2024; Ji et al., 2024b). This approach is feasible because most telecommunication optical fibers in urban areas are deployed along roadsides, making traffic the primary seismic source for interferometry. However, conventional ambient noise interferometry has limitations, including long cross-correlation periods required for convergence, diminished temporal resolution, and challenges in capturing high-frequency signals, which are often attenuated in ambient noise (Yuan et al., 2024). These issues are exacerbated by the heterogeneous and localized nature of noise sources in urban areas, which contrasts with the assumption of uniformly distributed ambient sources typically required for ambient noise interferometry. In particular, roadway vehicles have distinct characteristics, such as vehicle weight, size, and driving speed, that impact the induced surface waves and the inversion of near-surface structures. Directly analyzing surface wave signals from vehicles of varying weights and speeds may result in low-quality near-surface imaging. For instance, we observed that picked fundamental mode dispersion curves from vehicle-induced DAS signals are robust at frequencies above 5 Hz but not at lower frequencies. This is because the dominant traffic type, small/light vehicle, primarily excites surface waves at frequencies above 5 Hz. Therefore, understanding the effects of vehicle characteristics on the induced surface waves is important for improving the quality of near-surface imaging using vehicle-induced DAS signals in urban environments.
@@
In this study, we characterize surface waves generated by moving vehicles of different weights/sizes (sizes are closely correlated with weights) and speeds in the Stanford DAS-2 experiment. Building on our recent work (Yuan et al., 2024), we first use a specialized Kalman filter algorithm (Liu et al., 2023) to track vehicle locations and then classify vehicles by the peak values of their average quasi-static signals into weight and speed categories : light-weight (smaller than the mode value; peak value $\leq$0.57), heavy-weight (laying on the distribution tail; peak value $>$1.2), and mid-weight (in between; 0.57 $<$ peak value $\leq$1.2),. Vehicles are also categorized by their arrival times at DAS into traveling speed categories – slow, mid-speed, and fast – using the mean speed plus or minus one standard deviation as thresholds. Vehicle-induced surface wave windows are then selected using the tracked vehicle trajectories. Virtual shot gathers (VSGs) for the same class of vehicles are retrieved from their surface wave windows and stacked to increase the signal-to-noise ratio (Yuan et al., 2024). A VSG is a shot gathered as if a seismic source were excited at a specific location. This allows us to analyze the phase velocity dispersion at the location and further provides insights into how vehicle characteristics influence the induced surface waves. We also quantify the uncertainties in the picked dispersion curves as the number of stacked VSGs increases by bootstrapping the standard deviation of the phase velocities. Our results show that stacking VSGs reduces the uncertainty of picked dispersion curves, with convergence occurring after signals from approximately 55 vehicles are stacked. Heavy-weight vehicles show higher signal-to-noise ratios and reduced uncertainty in phase velocity estimates, particularly at lower frequencies (e.g., <5 Hz), allowing for better constraints of near-surface structures at greater depths. We also find that differences in the picked dispersion curves for vehicles with different speeds are less pronounced than those due to vehicle weight. In particular, for fine-grained temporal monitoring (e.g., daily monitoring) of near-surface characteristics, the low-frequency phase velocities (e.g., $<$5 Hz) should be derived from stacked heavy-weight vehicle VSGs and the high-frequency velocities from stacked VSGs of all classes of vehicles until convergence is achieved; for coarse-grained monitoring (e.g., weekly monitoring), the best dispersion curve should come from stacked heavy-weight vehicle VSGs, which is more efficient and accurate.
In summary, our results suggest that judiciously selecting certain classes of vehicles can be beneficial in processing vehicle-induced surface waves. The proposed approach in this study offers three key advantages:
  1. 1.
    By stacking the results of automatically tracked vehicles, we achieve highly accurate VSGs that capture detailed subsurface information with a high temporal resolution of up to daily intervals, a level of resolution challenging to obtain with conventional ambient noise interferometry.
  2. 2.
    Targeting specific vehicle classes further improves the signal-to-noise ratio in VSGs and enhances dispersion analysis accuracy, especially at lower frequencies. This leads to improved phase velocity estimation and greater sensitivity for imaging deeper subsurface structures.
  3. 3.
    The computational cost is minimized by only cross-correlation to the surface wave windows generated by the selected vehicle classes, significantly accelerating the processing workflow for large-scale applications.
@@ 3.1 Surface wave window selection based on vehicle tracking and characterization
Additionally, vehicle weights are estimated using the quasi-static component of the DAS signals. Assuming linear elasticity of the road and near-surface structures, the prominence amplitude of the quasi-static signal is approximately proportional to the vehicle’s weight. Heavier vehicles produce larger prominence amplitudes. Therefore, vehicles’ relative weights are estimated by computing the average of the quasi-static signal prominence, $\frac{1}{K}\sum_{k=1}^{K}P_{k}$, where $P_{k}$ is the prominence amplitude (or peak value) of the vehicle’s quasi-static signal at DAS channel $k$. The vehicle’s axle length and configuration can also be estimated from its quasi-static signal (Liu et al., 2023; Yuan et al., 2024). Yet, since vehicle length and weight are closely correlated, this study simplifies the analysis without considering different axle lengths and configurations.
@@ 5.1 Characterization of vehicle-induced DAS signal
We begin by characterizing the vehicle-induced DAS signals based on the estimated weights/sizes of the vehicles. To control for the variable of traveling speed, we select vehicles traveling within one standard deviation of the mode speed of all vehicles. We extract the quasi-static signals for each DAS channel by isolating a time window from the band-passed DAS signal (0.1–1 Hz), centered on the vehicle’s arrival time at that channel. Figure 2a displays the histogram of peak values of the average quasi-static signals of these vehicles. The peak value is defined as the average prominence amplitude of the vehicle’s quasi-static signal across all DAS channels within each surface wave window. We categorize the vehicles into three classes based on the peak values of their average quasi-static signals: vehicles with peak values smaller than the mode (which equals 0.57) are classified as light; those with peak values laying on the distribution tail (peak value $>1.2$) are classified as heavy; and those with values in between these thresholds are considered mid-weight. During the half-month period, we recorded 734 light, 1,058 mid-weight, and 103 heavy vehicles, respectively. Heavy vehicles, on average, weigh approximately seven times more than light vehicles. Figure 2b illustrates the average quasi-static DAS signals for vehicles having different weights, where the shaded areas highlight the signal ranges. Figure 2c shows the power spectral densities (PSD) of the surface wave signals band-passed between 1 and 30 Hz for different vehicle weights, with shaded areas representing the range of PSDs. We observe that heavier vehicles tend to produce both quasi-static and surface wave signals with stronger amplitudes. These signals also exhibit higher signal-to-noise ratio and greater surface wave propagation distances, as evidenced in the example signals (Figures 2d-f). Additionally, surface waves induced by heavy vehicles have a more prominent mode in the low-frequency range (2.5–5 Hz) compared to lighter ones.
@@
We also investigate the DAS signals induced by vehicles traveling at different velocities, as illustrated in Figure 3. Initially, to minimize the influence of vehicle weight and size, we select vehicles whose peak quasi-static signals are within one standard deviation of the mode peak value. While this approach reduces the influence of vehicle weight and ensures a sufficient sample size for analysis, it does not fully disentangle the complex interplay between vehicle weight and velocity. Future research could leverage a more controlled dataset comprising vehicles with similar weights but varying speeds to refine our understanding of how velocity impacts vehicle-induced DAS signals. Subsequently, these vehicles are categorized into slow, mid-speed, and fast groups based on their average speeds, separated using one standard deviation ($\mu_{s}\pm\sigma_{s}$), as shown in Figure 3a. There are 336 slow, 1,442 mid-speed, and 330 fast vehicles, respectively. Figures 3d–f show the signal examples from vehicles traveling at different speeds with their traveling trajectories indicated by dashed lines, where a smaller slope corresponds to a faster speed. Given the limited range of driving velocities, the surface wave signals among vehicles with varying speeds are not significantly different. Still, we note that surface waves induced by faster vehicles show slightly stronger amplitudes, which could be attributed to the greater kinetic energy and the corresponding larger impact forces exerted on rough road surfaces (Jazar and Marzbani, 2024). Further research could explore the potential for more energetic surface wave signals generated by vehicles traveling at higher velocities, such as those on freeways, where pronounced vehicle-road interactions are likely due to increased velocity and the road roughness (Yuan, 2023).
@@ 5.2 Surface wave dispersion analysis
We find that the quality of both the VSGs and dispersion measurements improves with increasing vehicle weight. The retrieved VSGs from heavier vehicles (Figure 4c) exhibit less noise in the negative time lag section, likely due to more accurate vehicle tracking and reduced ambient noise, as discussed in Yuan et al. (2024). Heavier vehicles generate more pronounced quasi-static signals, facilitating more accurate vehicle detection and tracking, as well as higher signal-to-noise ratio surface wave windows. The stronger signals from the tracked heavy vehicles effectively overshadow the low-amplitude anti-causal signals in the VSGs, which are typically generated by vehicles traveling in distant opposing traffic lanes. This results in improved clarity and reliability of the VSGs and their corresponding dispersion spectra. On the dispersion spectra (Figures 4d–f), the fundamental mode of dispersion shows greater continuity as vehicle weight increases, especially at low frequencies (<5 Hz). From the picked dispersion curves for vehicles of different weights (Figures 4g–i), we observe that those from heavy vehicles are more concentrated and show less uncertainty than those from light and mid-weight vehicles. Thus, low-frequency dispersion curves, which resolve deeper structures, can be more reliably measured from heavy vehicles. Retrieving VSGs from heavy vehicles also improves the measurement of higher-order dispersion. Compared to the results from light vehicles (Figure 4d), the higher mode dispersion curves are more distinguishable for mid-weight and heavy vehicles (Figures 4e and f). Considering that light-to-middle-weight vehicles are approximately 17 times more than heavy vehicles during the same fixed time period, we retrieved VSGs by stacking 1,020 light-to-middle-weight vehicle signals and calculated the corresponding dispersion spectra for comparison, as shown in Figure S4. Despite stacking 1,020 VSGs, the resulting dispersion spectra for light-to-middle-weight vehicles still do not match the quality achieved with just 60 heavy vehicle VSGs. When compared with ambient noise interferometry,
@@ 5.3 Inversion of shear-wave velocity
Beyond analyzing the features of the vehicle-induced DAS signals and the dispersion spectra, we inverted the dispersion curves to probe the subsurface structures. We invert for parameters including thickness and shear-wave velocity ($V_{s}$) for each layer, assuming a six-layer model. The density of each layer is approximated using an empirical relationship with the P-wave velocity as in Ji et al. (2024a). From the VSGs, we estimate that the ratio between P-wave and S-wave velocities ranges from 2 to 4 at this location by measuring their travel time. Based on the relationship between P-wave velocity, S-wave velocity, and Poisson’s ratio, we calculated a Poisson’s ratio range of 0.33 to 0.47, which was used in the inversion. The parameter ranges for inversion are specified in Table LABEL:tab:param.
@@ 6.1 Uncertainty quantification of dispersion curves
We gradually increase $N$, the number of VSGs stacked, to study the convergence of the uncertainty of the fundamental mode dispersion curves (2.5 to 14 Hz) and to estimate how many vehicles are needed to achieve convergence in dispersion curve picking. For each bootstrapping, M is set to 30 to ensure sufficient samples for calculating the standard deviation while maintaining manageable data volumes for processing. We apply the bootstrapping method at both pivot locations of 600 and 700 meters. Figures 8a and d show that the uncertainties (standard deviation) converged after stacking approximately 55 VSGs across all cases. The 600-meter pivot location shows a similar result, as detailed in the Supplemental Materials (Figure S3). The converged standard deviation of dispersion curves for heavy vehicles is about four times smaller than that for light and mid-weight vehicles at the 700-meter pivot location and five times smaller at the 600-meter location. The converged standard deviations of dispersion curves for vehicles with different speeds are similar. However, stacking VSGs from fewer slow-moving vehicles (e.g., fewer than 10 vehicles) results in larger uncertainty in dispersion curves. We also sampled an equal number of surface wave windows from each vehicle class to create subsets with the same number of observations. Subsequently, we perform an additional uncertainty quantification of the picked dispersion curves based on the percentage of vehicles relative to the total number of vehicles within each subset, as shown in Figures 8b and e. In particular, we calculate the standard deviation of the picked dispersion curves while incrementally increasing the percentage of VSGs stacked. This approach ensures a fair comparison of uncertainty, accounting for the differing numbers of recorded vehicles across weight classes during the fixed data collection period. Consistent with previous observations, stacking the same percentage of vehicles reveals that the low-frequency dispersion curves for heavy vehicles exhibit lower uncertainty (standard deviation) compared to those for light and mid-weight vehicles. The percentage of stacked vehicles is capped at 50% due to limitations in the bootstrapping algorithm. The algorithm estimates uncertainty by randomly sampling vehicle-induced surface wave windows with replacement from all available windows within the same vehicle class, repeating this process multiple times. A very large sampling size introduces a significant issue: the limited population size leads to a high probability of repeatedly sampling the same vehicles. This reduces the variability and independence of the subsamples, undermining the statistical robustness and accuracy of the bootstrapped estimates.
@@
Figure 8: Uncertainty quantification of the picked dispersion curves for different numbers of vehicles with varying weights and speeds. The uncertainties are quantified by bootstrapping the standard deviations of the estimated fundamental mode dispersion curves. (a) shows the standard deviations of the fundamental mode dispersion curves for vehicles with different weights at pivot locations of 700 meters. (b) shows the standard deviations of the fundamental mode dispersion curves based on the percentage of vehicles’ VSGs stacked with varying vehicle weights. (c) displays the converged standard deviations of the fundamental mode dispersion curves at various frequencies for vehicles with different weights. (d), (e) and (f) present the same data as (a), (b), and (c) but for vehicles with different speeds. The color version of this figure is available only in the electronic edition.
@@
We further investigate the converged uncertainty across different frequencies. Figure 8c and d display the converged standard deviations of the fundamental mode phase velocities at various frequencies for different vehicle classes. At frequencies below 5 Hz, the converged uncertainties for heavy vehicles are significantly smaller than those for light and mid-weight vehicles. For example, at 3 Hz, the converged uncertainty for heavy vehicles is approximately 3$\times$ smaller than that for light and mid-weight vehicles at the 700-meter pivot location. Another pivot location at 600 meters presents similar results, as shown in the Supplemental Materials (Figure S3). At higher frequencies, the converged uncertainties of phase velocities for vehicles of different weights are similar. Moreover, the converged uncertainties of phase velocities at different frequencies remain consistent across vehicles traveling at different speeds.